# Uncertainty-Weighted Image-Event Multimodal Fusion for Video Anomaly Detection

### v1

@@ Abstract
Most existing video anomaly detectors rely solely on RGB frames, which lack the temporal resolution needed to capture abrupt or transient motion cues—key indicators of anomalous events. To address this limitation, we propose Image-Event Fusion for Video Anomaly Detection (IEF-VAD), a framework that synthesizes event representations directly from RGB videos and fuses them with image features through a principled, uncertainty-aware process. The system (i) models heavy-tailed sensor noise with a Student’s-t likelihood, deriving value-level inverse-variance weights via a Laplace approximation; (ii) applies Kalman-style frame-wise updates to balance modalities over time; and (iii) iteratively refines the fused latent state to erase residual cross-modal noise. Without any dedicated event sensor or frame-level labels, IEF-VAD sets a new state of the art across multiple real-world anomaly detection benchmarks. These findings highlight the utility of synthetic event representations in emphasizing motion cues that are often underrepresented in RGB frames, enabling accurate and robust video understanding across diverse applications without requiring dedicated event sensors. Code and models are available at https://github.com/EavnJeong/IEF-VAD.
@@ 5 Conclusion
IEF-VAD proves that motion-centric synthetic event streams distilled from ordinary videos can be fused with RGB features to push video-anomaly detection beyond the limits of frame-based models. The approach establishes new state-of-the-art scores on UCF-Crime, XD-Violence, ShanghaiTech, and MSAD, while a value-level masking study shows its precision weights shift adaptively across latent dimensions—making the fusion logic transparent and reliable. We expect these findings to catalyse broader use of synthetic event data and to drive future work on uncertainty-aware multimodal learning for real-time surveillance.
Limitations and Future Work. While our study advances uncertainty-guided fusion, several limitations invite further investigation. First, approximating each modality’s noise with a diagonal covariance neglects cross-feature correlations; future work could adopt lightweight structured or low-rank covariance estimators to enrich value-level fusion. Second, the fixed regularization weights $\lambda_{1,2}$ and Student-t degrees-of-freedom $\nu$ limit adaptability; jointly learning or meta-learning these parameters would enable the model to recalibrate trust as modality quality shifts. Third, replacing simple weight averaging with uncertainty-conditioned cross-modal attention, hierarchical gating, or iterative message passing may yield a more expressive fusion mechanism that leverages inter-modal disagreement as informative signal, paving the way for stronger anomaly detection and broader multimodal perception capabilities.

### v2

@@ Abstract
Most existing video anomaly detectors rely solely on RGB frames, which lack the temporal resolution needed to capture abrupt or transient motion cues—key indicators of anomalous events. To address this limitation, we propose Image-Event Fusion for Video Anomaly Detection (IEF-VAD), a framework that synthesizes event directly from RGB videos and fuses them through a principled, uncertainty-aware process. The system (i) models heavy-tailed sensor noise with a Student’s t likelihood, deriving value-level inverse-variance weights via a Laplace approximation; (ii) applies Kalman-style frame-wise updates to recalibrate modality weights based on temporal context; and (iii) iteratively refines the fused latent state to remove residual cross-modal noise. Without any dedicated event sensor or frame-level labels, IEF-VAD sets a new state of the art across multiple real-world anomaly detection benchmarks. These findings highlight the utility of synthetic event in emphasizing motion cues that are often underrepresented in RGB frames, enabling accurate and robust video understanding across diverse applications without requiring dedicated event sensors. The code is available at: https://github.com/EavnJeong/IEF-VAD
@@ 5 Conclusion
IEF-VAD proves that motion-centric synthetic event streams distilled from ordinary videos can be fused with RGB frame to push video-anomaly detection beyond the limits of frame-based models. The approach establishes new state-of-the-art scores on UCF-Crime, XD-Violence, ShanghaiTech, and MSAD, while a value-level masking study shows its weights shift adaptively across latent dimensions, revealing a reliable reliance on modality-specific cues. We expect these findings to catalyse broader use of synthetic event data and demonstrate its potential for principled, effective multimodal fusion in video understanding.
Limitations and Future Work. While our study advances uncertainty-guided fusion, several limitations invite further investigation. First, approximating each modality’s noise with a diagonal covariance neglects cross-feature correlations; future work could adopt lightweight structured or low-rank covariance estimators to enrich value-level fusion. Second, the fixed regularization weights $\lambda_{1,2}$ and Student-t degrees-of-freedom $\nu$ limit adaptability; learning these parameters jointly with the model, or via meta-learning, would allow the system to adjust its reliance on each modality based on their varying quality. Third, replacing simple weight averaging with uncertainty-conditioned cross-modal attention, hierarchical gating, or iterative message passing may yield a more expressive fusion mechanism that leverages inter-modal disagreement as informative signal, paving the way for stronger anomaly detection and broader multimodal perception capabilities.